# TRIPS: Trilinear Point Splatting for Real-Time Radiance Field Rendering

### v1

@@ 1 Introduction
A common strategy to handle these artifacts is to use scene-specific optimization methods, known as inverse rendering. This allows for the adjustment of the scene’s texture, geometry, and camera parameters to align the rendering with the photograph. Prominent techniques in this domain incorporate per-point descriptors [RFS22, $*{}^{*}$20, KPLD21], explicit optimization of point sizes via Gaussians [KKLD23] and learned neural refinement networks [TZN19, RFS22, $*{}^{*}$22]. While this generally extends render times, it significantly enhances visual quality.
@@
  • An implementation of the method resulting in high-quality real-time renderings under varying capturing conditions at https://github.com/lfranke/TRIPS
Figure 1: Our pipeline: TRIPS renders a point cloud trilinearly as 2$\times$2$\times$2 splats into multi-layered feature maps with the results being passed though our small neural network, containing only a single gated convolution per layer. This pipeline is completely differentiable, so that point descriptors (colors) and positions, as well as camera parameters are optimized via gradient descent.
@@ 2.1 Novel View Synthesis and Traditional Approaches
Traditionally, real-world novel view synthesis relies on image-based rendering techniques. Commonly, Structure-from-Motion (SfM) techniques [SSS06, SF16] allow camera parameter estimations from a set of photographs which are then used for directly warping source image colors to a target view [DYB98, CDSHD13]. This relies on accurate proxy geometry (usually point clouds or meshed), commonly enhanced via Multi-View Stereo (MVS) [SZPF16, $*{}^{*}$07]. In real world datasets however, these techniques can present camera miscalibrations and erroneous geometry [SK00]. For image-based rendering, this can lead to warping artifacts, especially near object boundaries, or can cause blurring of details. Recently, pipelines enhanced by neural rendering [$*{}^{*}$22] provided powerful tools to lessen these artifacts.
@@ 2.2 Neural Rendering and Scene Representations
In the last years, multiple variants of deep learning for novel view synthesis were introduced. Within proxy-based pipelines, several works have replaced the blending operation by a deep neural networks [RK21, $*{}^{*}$18, RK20] or learned textures [TZN19] during the warping stage. Other approaches use multi-plane images [$*{}^{*}$19, $*{}^{*}$19, TS20, $*{}^{*}$18] or estimate a warping fields [FNPS16, GKSL16, $*{}^{*}$16] to avoid the need of an scene specific proxy geometry.
This led the way for volumetric scene representations [PZ17] enhanced with deep learning [$*{}^{*}$20, $*{}^{*}$19] and rendered via ray-marching. Neural Radiance Fields (NeRFs) [$*{}^{*}$21] furthermore showed that compressing a full 3D scene into a Multilayer Perceptron (MLP) achieve great results in this regard. This representation however is challenging in its own right, which follow-up works improve upon: long training times [$*{}^{*}$21, CBLPM21, MESK22, $*{}^{*}$21, TRS22], many well distributed input views [CBLPM21, YYTK21, KD23] and rendering times [MESK22, $*{}^{*}$21, $*{}^{*}$21]. Improvements in quality [$*{}^{*}$23, $*{}^{*}$21] allow NeRFs to surpass visual quality of many proxy-based approaches, however render times are still challenging, e.g. MipNeRF-360 [$*{}^{*}$22] ranging in the order of seconds per image and training needing dozens of hours.
Lately, discretizing parts of the scene space [$*{}^{*}$21, $*{}^{*}$21] or even replacing parts of it via voxel grids [$*{}^{*}$22], octrees [$*{}^{*}$22] or tensor factorization [$*{}^{*}$22] shrink computational costs as MLPs can be smaller or even removed. In this area, InstantNGP [MESK22] made waves as it uses hash-grids and a highly optimized MLP implementation for faster rendering and trainings speeds while retaining many qualitative advantages of NeRFs.
@@ 2.3 Real-Time Rendering for Radiance Fields via Points
In the domain of real-time radiance field rendering, point clouds as an explicit proxy representation remain a great option. Point clouds are easily captured via LiDAR-based mapping [LXG22], RGB-D cameras with fusion techniques [$*{}^{*}$17, $*{}^{*}$16, $*{}^{*}$13] and SfM/MVS techniques [SZPF16]. They represent a unstructured set of samples in space, with varying distances to neighbors, but true to the originally captured data. Rendering these can be very fast [SKW21, SKW22, SKW19], and augmenting points with neural descriptors [RFS22, $*{}^{*}$20, RALB22] or optimized attributes [KPLD21, $*{}^{*}$22] provide high quality renderings using differentiable point renderers [WGSJ20, $*{}^{*}$19]. However, discrete rasterization of points can cause aliasing [SKW22] or overdraw [RFS22] if many points are rendered to the same pixel.
@@
In world-space hole-filling, points are represented as oriented discs, often termed "splats" or "surfels", with disc radii precomputed based on point cloud density. To reduce artifacts between neighboring splats, these discs can be rendered using Gaussian alpha-masks and combined with a normalizing blend function [$*{}^{*}$04, PZVBG00, ZPVBG01]. Recent techniques optimize splat sizes [KKLD23, ZBRH22] or improve quality with neural networks [$*{}^{*}$20]. For performance, overdraw poses a major issue as splats tend to overlap a lot. Thus, special care has to be taken regarding the amount of splats drawn. 3D Gaussian Splatting [KKLD23] can be considered state of the art in this domain. They combine anisotropic Gaussians with a very fast tiled renderer and optimize splat sizes via gradient descent. However, limiting Gaussian numbers is necessary to avoid performance hits, which in turn can lead to over-blurred small detailed elements.
The second direction involves screen-space hole-filling, where points, often rendered as tiny splats, are post-processed either through traditional methods [PGA11, MKC07, GD98] or using convolutional neural networks (CNNs) [$*{}^{*}$20, $*{}^{*}$19, SCCL20]. While these techniques bridge large point distances, their need for a large receptive field can result in artifacts or performance issues. A multi-resolution pyramid rendering approach mitigates this by assigning different network layers to varied $*{}^{*}$20, RALB22], albeit reintroducing overdraw issues at lower layers [RFS22]. Notably, ADOP [RFS22] excels in screen-space hole-filling, enabling the rendering of hundreds of millions of points for sharp object visualization [SKW22], but encounters challenges with temporal aliasing and substantial hole-filling.
@@ 3 Method
Fig. 1 provides an overview of our rendering pipeline. The input data consists of images with camera parameters and a dense point cloud, which can be obtained through methods like multi-view stereo [SZPF16] or LiDAR sensing. To render a specific view, we project the neural color descriptors of each point into an image pyramid using the TRIPS technique (as detailed in Sec. 3.1) and blends them (Sec. 3.2). Subsequently, a compact neural reconstruction network (described in Sec. 3.3) integrates the layered representation, followed by the application of a spherical harmonics module (discussed in Sec. 3.4) and a tone mapper that transforms the resulting features into RGB colors.
@@
$\text{I}=\Phi(C,R,t,x,E,\upsilon,\tau,\gamma),$ (1)
where $C$ are the camera intrinsics, $(R,t)$ the extrinsic pose of the target view, $x$ the position of the points, $E$ the optional environment map, $\upsilon$ the world space size of the points, $\tau$ the neural point descriptors and $\gamma$ the transparency for each point.
@@ 3.1 Differentiable Trilinear Point Splatting
Using $C$ and $(R,t)$, we project each point position $(x_{w},y_{w},z_{w})$ to continuous (non-rounded) screen space coordinates $(x,y,z)$ and each world-space point size $s_{w}$ to screen space size $s$ with the camera’s focal length $f$:
@@
Next, we render these points as a $2\times 2\times 2$ splats bilinearly and handle point size by splatting into two neighboring resolution layers $L$ with $L_{lower}=\lfloor\log(s)\rfloor$ and $L_{upper}=\lceil\log(s)\rceil$, as shown in Fig. 2. The final opacity value $\gamma$ that is written to the image pyramid for pixel $(x_{i},y_{i},s_{i})$ with $s_{i}=\lfloor 2^{L}\rfloor$ is
@@
$\iota$ =$=\begin{cases}1-|s-s_{i}|&s\geq 1\\ \epsilon+(1-\epsilon)s&s_{i}=0\land s<1\end{cases}$ (5)
@@ 3.2 Multi Resolution Alpha Blending
Figure 4: Our design of one gated convolution block that processes the features of the image pyramid.
@@ 3.3 Neural Network
The result produced by our renderer consists of a feature image pyramid comprising $n$ layers. These individual layers are finally consolidated into a single full-resolution image by a compact neural network, as depicted in Fig. 1. Our network architecture incorporates a single gated convolution [$*{}^{*}$19] in each layer, featuring a self-bypass connection and a filter size of 32. Additionally, we include a bilinear upsampling operation for all layers except the final one, merging the output with the subsequent level. This configuration is shown in Fig. 4 and resembles an efficient decoder network, due to its restrained number of features, pixels, and convolutional operations.
Unlike well-established hole-filling neural networks [$*{}^{*}$20, RFS22, RALB22], our approach demands a significantly smaller and more efficient network. This reduced network size stems from the fact that our renderer is adept at filling gaps autonomously and generates smooth output. Consequently, the network’s primary task is to learn minimal hole-filling and outlier removal, allowing it to concentrate its efforts on high-quality texture reconstruction.
@@ 3.4 Spherical Harmonics Module and Tone Mapping
To model view dependent effects and camera-specific capturing parameters (like exposure time), we optionally interpret the network output as spherical harmonics (SH) coefficients, convert them to RGB colors, and finally pass the result to a physically-based tone mapper. This allows the system to make use of explicit view directions. The SH-module makes use of 3 bands, which corresponds to 27 input coefficients. The tone mapper follows the work of Rückert et al. [RFS22], which models exposure time, white balance, sensor response, and vignetting.
@@ 3.6 Implementation Details
The renderer is implemented in three stages: collecting, splatting and accumulation, albeit diverging from other state-of-the-art multi-layer blending strategies, this turned out to work best in our scenario [FHSS18, LZ21, VVP20]. We first project each point $p_{i}$ to the desired view and collect each point’s $(x,y,z)$ as well as $s$ in a buffer, and also count how many elements are mapped to each pixel. This counting is then used for an offset scan to allocate continuous arrays for all layers combined. The following splatting pass duplicates each point and stores a pair of $(z,i)$ in each pixels’ list.
@@
InstantNGP 0.309 21.74 0.692 0.289 25.94 0.697 0.539 15.34 0.455 0.316 13.45 0.803
Mip-NeRF 360 0.203 24.61 0.789 0.174 28.23 0.796 0.676 12.20 0.357 0.322 15.21 0.833
Gaussian Spl. 0.197 24.63 0.818 0.166 26.94 0.792 0.467 15.30 0.470 0.184 18.77 0.878
ADOP 0.121 23.78 0.802 0.163 23.26 0.707 0.179 20.49 0.650 0.117 21.47 0.899
Ours 0.099 24.64 0.808 0.122 25.94 0.772 0.179 20.38 0.633 0.107 21.36 0.887
@@ 4.1 Setup and Datasets
We have evaluated our approach on several scenes from the Tanks&Temples [KPZK17] and the MipNeRF-360 [$*{}^{*}$22] datasets. Additionally, we use the Boat and Office scene from Rückert et al. [RFS22] to evaluate robustness towards difficult input conditions. The former contains outdoor auto-exposed images while the later is a office floor with multiple distinct room and a large LiDAR point cloud, but sparsely placed cameras.
From Tanks&Temples, we use the intermediate set containing eight scenes: Train, Playground, M60, Lighthouse, Family, Francis, Horse and Panther. These scenes are outdoor scenes captured under varying lighting conditions but with good spatial coverage and can be seen as a good baseline for robustness. The MipNeRF-360 dataset [$*{}^{*}$22] consists of 5 outdoor and 4 indoor scenes. This dataset was captured with controlled setups and has capture positions well suited for volumetric rendering with a hemispherical setup [KD23].
For the quantitative evaluation we use the LPIPS [$*{}^{*}$18], PSNR, and SSIM metrics. We note however, that neither of these metrics always reflect visual impression. Some approaches are trained with MSE-loss or SSIM and therefore naturally perform better in PSNR and SSIM. Our approach, on the other hand, is trained with VGG-loss and thus usually shows better scores on LPIPS. For a fair comparison, we recommend to look at all metrics and closely inspect the provided image and video comparison.

### v2

@@ 1 Introduction
A common strategy to handle these artifacts is to use scene-specific optimization methods, known as inverse rendering. This allows for the adjustment of the scene’s texture, geometry, and camera parameters to align the rendering with the photograph. Prominent techniques in this domain incorporate per-point descriptors [RFS22, ASK20, KPLD21], explicit optimization of point sizes via Gaussians [KKLD23] and learned neural refinement networks [TZN19, RFS22, KLR22]. While this generally extends render times, it significantly enhances visual quality.
@@
  • An implementation of the method resulting in high-quality real-time renderings under varying capturing conditions at:
    https://github.com/lfranke/TRIPS
Figure 1: Our pipeline: TRIPS renders and blends a point cloud trilinearly as 2$\times$2$\times$2 splats into multi-layered feature maps with the results being passed though our small neural network, containing only a single gated convolution per layer. Following, an optional spherical harmonics module and tone-mapper is used to produce the final image. This pipeline is completely differentiable, so that point descriptors (colors) and positions, as well as camera parameters are optimized via gradient descent.
@@ 2.1 Novel View Synthesis and Traditional Approaches
Traditionally, real-world novel view synthesis relies on image-based rendering techniques. Commonly, Structure-from-Motion (SfM) techniques [SSS06, SF16] allow camera parameter estimations from a set of photographs which are then used for directly warping source image colors to a target view [DYB98, CDSHD13]. This relies on accurate proxy geometry (usually point clouds or meshed), commonly enhanced via Multi-View Stereo (MVS) [SZPF16, GSC07]. In real world datasets however, these techniques can present camera miscalibrations and erroneous geometry [SK00]. For image-based rendering, this can lead to warping artifacts, especially near object boundaries, or can cause blurring of details. Recently, pipelines enhanced by neural rendering [TTM22] provided powerful tools to lessen these artifacts.
@@ 2.2 Neural Rendering and Scene Representations
In the last years, multiple variants of deep learning for novel view synthesis were introduced. Within proxy-based pipelines, several works have replaced the blending operation by a deep neural networks [RK21, HPP18, RK20, FRF23a] or learned textures [TZN19] during the warping stage. Other approaches use multi-plane images [MSOC19, STB19, TS20, ZTF18] or estimate a warping fields [FNPS16, GKSL16, ZTS16] to avoid the need of an scene specific proxy geometry.
This led the way for volumetric scene representations [PZ17] enhanced with deep learning [SMB20, STH19] and rendered via ray-marching. Neural Radiance Fields (NeRFs) [MST21] furthermore showed that compressing a full 3D scene into a Multilayer Perceptron (MLP) achieve great results in this regard. This representation however is challenging in its own right, which follow-up works improve upon: long training times [CXZ21, CBLPM21, MESK22, TMW21, TRS22], many well distributed input views [CBLPM21, YYTK21, KD23] and rendering times [MESK22, BMT21, NSP21]. Improvements in quality [BMV23, MBRS21] allow NeRFs to surpass visual quality of many proxy-based approaches, however render times are still challenging, e.g. MipNeRF-360 [BMV22] ranging in the order of seconds per image and training needing dozens of hours.
Lately, discretizing parts of the scene space [YLT21, HSM21] or even replacing parts of it via voxel grids [FKYT22], octrees [RWL22] or tensor factorization [CXG22] shrink computational costs as MLPs can be smaller or even removed. In this area, InstantNGP [MESK22] made waves as it uses hash-grids and a highly optimized MLP implementation for faster rendering and trainings speeds while retaining many qualitative advantages of NeRFs.
@@ 2.3 Real-Time Rendering for Radiance Fields via Points
In the domain of real-time radiance field rendering, point clouds as an explicit proxy representation remain a great option. Point clouds are easily captured via LiDAR-based mapping [LXG22], RGB-D cameras with fusion techniques [DNZ17, WSMG16, KLL13] and SfM/MVS techniques [SZPF16]. They represent a unstructured set of samples in space, with varying distances to neighbors, but true to the originally captured data. Rendering these can be very fast [SKW21, SKW22, SKW19], and augmenting points with neural descriptors [RFS22, ASK20, RALB22, FRF23b, HKT23] or optimized attributes [KPLD21, KLR22] provide high quality renderings using differentiable point renderers [WGSJ20, YSW19] or neural ray-based renderers [XXP22, OLN22, ACDS24]. However, discrete rasterization of points can cause aliasing [SKW22] or overdraw [RFS22] if many points are rendered to the same pixel.
@@
In world-space hole-filling, points are represented as oriented discs, often termed "splats" or "surfels", with disc radii precomputed based on point cloud density. To reduce artifacts between neighboring splats, these discs can be rendered using Gaussian alpha-masks and combined with a normalizing blend function [AGP04, PZVBG00, ZPVBG01]. Recent techniques optimize splat sizes [KKLD23, ZBRH22] or improve quality with neural networks [YCA20]. For performance, overdraw poses a major issue as splats tend to overlap a lot. Thus, special care has to be taken regarding the amount of splats drawn. 3D Gaussian Splatting [KKLD23] can be considered state of the art in this domain. They combine anisotropic Gaussians with a very fast tiled renderer and optimize splat sizes via gradient descent. However, limiting Gaussian numbers is necessary to avoid performance hits, which in turn can lead to over-blurring of small detailed elements.
The second direction involves screen-space hole-filling, where points, often rendered as tiny splats, are post-processed either through traditional methods [PGA11, MKC07, GD98] or using convolutional neural networks (CNNs) [ASK20, MGK19, SCCL20]. While these techniques bridge large point distances, their need for a large receptive field can result in artifacts or performance issues. A multi-resolution pyramid rendering approach mitigates this by assigning different network layers to varied resolutions[ASK20, RALB22, HFF23], albeit reintroducing overdraw issues at lower layers [RFS22]. Notably, ADOP [RFS22] excels in screen-space hole-filling, enabling the rendering of hundreds of millions of points for sharp object visualization [SKW22], but encounters challenges with temporal aliasing and substantial hole-filling.
@@ 3 Method
Fig. 1 provides an overview of our rendering pipeline. The input data consists of images with camera parameters and a dense point cloud, which can be obtained through methods like multi-view stereo [SZPF16] or LiDAR sensing. To render a specific view, we project the neural color descriptors of each point into an image pyramid using the TRIPS technique (as detailed in Sec. 3.1) and blend them (Sec. 3.2). Subsequently, a compact neural reconstruction network (described in Sec. 3.3) integrates the layered representation, followed by the application of a spherical harmonics module (discussed in Sec. 3.4) and a tone mapper that transforms the resulting features into RGB colors.
@@
$\text{I}=\Phi(C,R,t,x,E,s_{w},\tau,\alpha),$ (1)
where $C$ are the camera intrinsics, $(R,t)$ the extrinsic pose of the target view, $x$ the position of the points, $E$ the optional environment map, $s_{w}$ the world space size of the points, $\tau$ the neural point descriptors and $\alpha$ the transparency for each point.
@@ 3.1 Differentiable Trilinear Point Splatting
Using camera intrinsics $C$ and pose $(R,t)$, we project each point position $(x_{w},y_{w},z_{w})$ to continuous (non-rounded) screen space coordinates $(x,y,z)$ and each world-space point size $s_{w}$ to screen space size $s$ with the camera’s focal length $f$:
@@
Next, we render these points as a $2\times 2\times 2$ splats bilinearly and handle point size by splatting into two neighboring resolution layers $L$, as shown in Fig. 2. The resolution layers are selected to be the two closest in sizes to the projected size of the point with $L_{lower}=\lfloor\log(s)\rfloor$ and $L_{upper}=\lceil\log(s)\rceil$.
For each of the then selected eight pixels, we compute the contribution of the point to that pixel and augment its own transparency value value with it. The final opacity value $\gamma$ that is written to the image pyramid for pixel $(x_{i},y_{i},s_{i})$ with $s_{i}=2^{L}$ is
@@
$\iota$ =$=\begin{cases}1-\frac{|s-s_{i}|}{2^{L_{upper}}-2^{L_{lower}}}&s\geq 1\\ \epsilon+(1-\epsilon)s&s_{i}=0\land s<1\end{cases}$ (5)
@@ 3.2 Multi Resolution Alpha Blending
Figure 4: Our design of one gated convolution block that processes the features of the image pyramid with the number of channels passed through indicated at each step.
@@ 3.3 Neural Network
The result produced by our renderer consists of a feature image pyramid comprising $n$ layers. These individual layers are finally consolidated into a single full-resolution image by a compact neural network, as depicted in Fig. 1. Our network architecture incorporates a single gated convolution [YLY19] in each layer with a self-bypass connection and a feature size of 32. Additionally, we include a bilinear upsampling operation for all layers except the final one, merging the output with the subsequent level. This configuration is shown in Fig. 4 and resembles an efficient decoder network, due to its restrained number of features, pixels, and convolutional operations.
Unlike well-established hole-filling neural networks [ASK20, RFS22, RALB22], our approach demands a significantly smaller and more efficient network. This reduced network size stems from the fact that our renderer is adept at filling gaps autonomously and generates smooth output through trilinearly splatting points. Consequently, the network’s primary task is to learn minimal hole-filling and outlier removal, allowing it to concentrate its efforts on high-quality texture reconstruction.
@@ 3.4 Spherical Harmonics Module and Tone Mapping
To model view dependent effects and camera-specific capturing parameters (like exposure time), we optionally interpret the network output as spherical harmonics (SH) coefficients, convert them to RGB colors, and finally pass the result to a physically-based tone mapper. This allows the system to make use of explicit view directions. The SH-module makes use of spherical harmonics with degree 2, which corresponds to 27 input coefficients (9 coefficients per color channel). These coefficients are the output of the last convolution of our network. The tone mapper follows the work of Rückert et al. [RFS22], which models exposure time, white balance, sensor response, and vignetting.
@@ 3.6 Implementation Details
The renderer is implemented in three stages: collecting, splatting and accumulation, albeit diverging from other state-of-the-art multi-layer blending strategies, this turned out to work best in our scenario [FHSS18, LZ21, VVP20]. We first project each point $(x_{w},y_{w},z_{w})$ to the desired view and collect each point’s $(x,y,z)$ as well as point size $s$ in a buffer, and also count how many elements are mapped to each pixel. This counting is then used for an offset scan to index into one continuous arrays for all layers. The following splatting pass duplicates each point and stores a pair of $(z,i)$ (with $i$ an index to the stored information) in each pixels’ list.
@@
In contrast to Kerbl and Kopanas et al. [KKLD23], we use this per-pixel sorting, which proved to be faster for us then global sorting. This is mostly due to the higher amount and smaller sizes of points in our approach.
@@
InstantNGP 0.475 21.74 0.692 0.374 25.94 0.697 0.598 15.34 0.455 0.544 13.45 0.801
Mip-NeRF 360 0.340 24.61 0.789 0.286 28.23 0.796 0.680 12.20 0.357 0.526 15.22 0.832
Gaussian Spl. 0.300 24.63 0.818 0.278 26.94 0.792 0.544 15.30 0.470 0.371 18.77 0.878
ADOP 0.229 23.78 0.802 0.285 23.26 0.707 0.301 20.49 0.650 0.279 21.47 0.899
Ours 0.213 24.64 0.808 0.233 25.94 0.772 0.301 20.38 0.633 0.271 21.36 0.887
@@ 4.1 Setup and Datasets
We have evaluated our approach on several scenes from the Tanks&Temples [KPZK17] and the MipNeRF-360 [BMV22] datasets. Additionally, we use the Boat and Office scene from Rückert et al. [RFS22] to evaluate robustness towards difficult input conditions. The former contains outdoor auto-exposed images while the later is a office floor with multiple distinct room and a large LiDAR point cloud, but sparsely placed cameras.
From Tanks&Temples, we use the intermediate set containing eight scenes: Train, Playground, M60, Lighthouse, Family, Francis, Horse and Panther. These scenes are outdoor scenes captured under varying lighting conditions but with good spatial coverage and can be seen as a good baseline for robustness. The MipNeRF-360 dataset [BMV22] consists of 5 outdoor and 4 indoor scenes. This dataset was captured with controlled setups and has capture positions well suited for volumetric rendering with a hemispherical setup [KD23]. We use half resolution for images of this dataset, resulting in resolutions of around $2500\times 1600$ px for outdoor and $1550\times 1030$ px for indoor scenes. For results with the resolutions used in related works (outdoor: quarter resolution; indoor half resolution), take a look at the Appendix, Tabs. 10-13.
Point clouds of all scenes were acquired via COLMAP’s MVS [SZPF16], except Office which was captured by LiDAR.
For the quantitative evaluation we use the LPIPSVGG [ZIE18], PSNR, and SSIM metrics. We note however, that neither of these metrics always reflect visual impression. Some approaches are trained with MSE-loss or SSIM and therefore naturally perform better in PSNR and SSIM. Our approach, on the other hand, is trained with VGG-loss and thus usually shows better scores on LPIPS. For a fair comparison, we recommend to look at all metrics and closely inspect the provided image and video comparison.
In all experiments, we leave every 8th view out for testing. This is the same train/test split as used in current related work [BMV22, KKLD23].